OpenStudy (anonymous):

Find the exact value. tan 75 degrees

OpenStudy (kamibug):

Do you have a scientific calculator? :)

OpenStudy (anonymous):

not with me sadly

jhonyy9 (jhonyy9):

calculi tan(45+30)=?

jhonyy9 (jhonyy9):

than you dont know it you can rewrite it like sin(45+30) ---------- = ? cos(45+30)

OpenStudy (anonymous):

\[\tan 75=\tan \left( 45+30 \right)=\frac{ 1+\tan 30 }{ 1-\tan 30 }=?\]

OpenStudy (anonymous):

@jhonyy9 I still don't understand and do not have a calculator with me to solve this problem.

jhonyy9 (jhonyy9):

tan 30 = sin30/cos30 = 1/2/sqrt3/2 = 1/2 *2/sqrt3 = 1/sqrt3 = sqrt3 /3 so now just substitute in this formula what have wrote to you surjithayer above - ok ?